OpenStudy (anonymous):

sin(x+pi/4)=sin(x−pi/4)−1 my answer is 3pi/4 and 5pi/4. i don't think it's correct, can somebody help me please?

OpenStudy (mertsj):

I think I might try using the formulas for sin(x+y) and sin(x-y). Have you tried that?

OpenStudy (anonymous):

yeah, that's how i came up with those answers

OpenStudy (anonymous):

the sum and diference identities, right?

OpenStudy (mertsj):

\[\sin x \cos 45+\cos x \sin 45=\sin x \cos 45-\cos x \sin 45-1=\frac{\sqrt{2}}{2}(\sin x+\cos x)=\frac{\sqrt{2}}{2}(\sin x-\cos x)-1\]

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

umm...it's cut off

OpenStudy (mertsj):

So we have \[2\cos x \sin 45=1\]

OpenStudy (mertsj):

\[2(\frac{\sqrt{2}}{2})\cos x=1\]

OpenStudy (mertsj):

\[\cos x=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

wait, it's not = -1?

OpenStudy (anonymous):

i thought it was = -1, not 1

OpenStudy (mertsj):

\[x=\frac{3pi}{4}, \frac{5pi}{4}\]

OpenStudy (mertsj):

Yep. You are correct. negative 1

OpenStudy (mertsj):

Oh no. I had to go and make that mistake when satellite is watching.

OpenStudy (anonymous):

so there are two answers?

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i think there is also a way to write \[\sin(x+\frac{\pi}{4})-\sin(x-\frac{\pi}{4})\] as a single function of cosine

OpenStudy (mertsj):

yes. two answers between 0 and 2pi. Infinite number of solutions if you want them all.

OpenStudy (anonymous):

nvm that is what you did. i guess i mean another way but it amounts to the same thing

OpenStudy (anonymous):

yup, it's only in the interval of 0<x<2pi

OpenStudy (mertsj):

I think you are correct but who wanted to look that up.

OpenStudy (anonymous):

great, Thanks!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

yeah you can turn \(-\sin(x-\frac{\pi}{4})\) in to \(\cos(x+\frac{\pi}{4})\) then \[\sin(x+\frac{\pi}{4})+\cos(x+\frac{\pi}{4})=\sqrt{2}\cos(x)\] or something like that i always forget how to do these myininaya remembers